OpenStudy (anonymous):

2x + 5y = 15 5x + 10y = 35

OpenStudy (anonymous):

5y=15-2x y=3-(2/5)x 5x+10(3-(2/5)x) = 35 5x+30-4x = 35 x+30 = 35 x = 5 Plug in x into either formula and you get... 2(5)+5y = 15 10 + 5y = 15 5y = 5 y = 1 Therefore, x = 5, y = 1

OpenStudy (anonymous):

multiply first equation with 2 4x+10y=30............(1) 5x+10y=35.............(2) subtract 2nd equation from 1st one: -x=-5 x=5 now put the value of x in any one of the above equations...... 4(5)+10y=30 10y=30-20 10y=10 y=1

OpenStudy (anonymous):

thanks